Operations Report – February 06th

Crew 221 Operations Report 06Feb2022

SOL: 4

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 8pm)

Hours run: 13:00

From what time last night: 7:00pm

To what time this morning: 8:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 76%

Diesel Reading – 98%

Station Propane Reading – 61%

Ethanol Free Gasoline – gallons

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 355 gals

Static to Loft Pump used – yes

Water in GreenHab – 243 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 129

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 35%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: yes

Hours: 135

Beginning charge: 100%

Ending charge: 25%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: not used

Summary of Hab operations: nominal.

Summary of GreenHab operations: black corn experiment started, 4 trays were filled 3 with mars soil and one with regular soil for control, beans were plantes and watered with 2L. operational, watering of the plants carried out as per procedure

Summary of ScienceDome operations: 2 incubators runing, microbiology experiments going on and cultures are bieng grown in agar plates.

Summary of RAM operations: robot assembly is ongoing adn expected to continue.

Summary of any observatory issues: nominal

Summary of health and safety issues: no issues

Questions, concerns and requests to Mission Support:

Copyright © The Mars Society. All rights reserved. | Main Site